The length of a rectangular window is 5 feet more than its width, w. The area of the window is 36 square feet. Which equation could be used to find the dimensions of the window?
Thank you for your help.

Answer by CharlesG2(834)

"The length of a rectangular window is 5 feet more than its width, w. The area of the window is 36 square feet. Which equation could be used to find the dimensions of the window?
Thank you for your help."
area rectangle A = length l * width w = 36 square feet
A = (w + 5) * w = 36
A = w^2 + 5w = 36
w^2 + 5w - 36 = 0
(w + 9)(w - 4) = 0
w^2 -4w + 9w - 36 = 0 (use FOIL)
w^2 + 5w - 36 = 0
width is 4 feet, so length is 9 feet (since 4*9=36)
so whatever equation matches up with what is done here is the answer